OpenStudy (anonymous):

The four sides of a vegtable garden are 16 1/5 feet, 32 3/4 feet, 25 1/4 feet, 20 4/5 feet. How many feet of fencing are needed to go around the garden? First estimate the answer and then solve for the exact answer.

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

"fencing" usually mean around the outside - the perimeter. Add the four values. What's your estimate? Give it a rough shot. Seriously, just stare at the four numbers and say to yourself, "That's about...ummm...oh, I don't know 70." Something like that. Don't worry about how close it is. We're only worried about reasonableness. We're not actually trying to get the right answer, yet.

OpenStudy (anonymous):

@mathstudent55

OpenStudy (tkhunny):

So, you're not going to do it?

OpenStudy (anonymous):

you mean them all added up as one whole answer. I would say about 90

OpenStudy (tkhunny):

Perfect. That's a wonderful estimate for reasonableness. Knowing this, if we get 412, we will know immediately that something has gone horribly wrong. If we get -16, something will have gone wrong. The real answer should be somewhat int eh neighborhood of 90 - your reasonableness estimate. Okay, now how do we add them up exactly?

OpenStudy (anonymous):

93

OpenStudy (tkhunny):

Well, that doesn't really qualify as showing your work. First, check the units. Are the ALL in feet or will one or two require some conversion? It is VERY important to make sure your units are the same. If they are not, you must convert hem. I this case, they are all in feet. Good. 16 1/5 feet, 32 3/4 feet, 25 1/4 feet, 20 4/5 feet I would do it this way: (16 + 32 + 25 + 20) + (1/5 + 3/4 + 1/4 + 4/5) = 93 + 2 = 95. Okay, where did you go wrong when you managed on 93?

OpenStudy (anonymous):

i have no idea so 93 would be the estimated answer only

OpenStudy (tkhunny):

How did we get another estimate? I thought the estimate was 90. We were supposed to be working on the exact answer.

OpenStudy (anonymous):

oh snap sorry i got lost ok i get it so 90 is estimated answer and 95 is exact nswer